If 43 m of nichrome wire is to have a resistance of 14.0 ? at 20
Answer #1

for a nichrome wire at 20 C

Resistivity ? = 1 * 10^-6 ohm . m

then

As

R = ? * L/A

14 = 1 * 10^-6 * 43/A

A = 3.07 * 10^-6 m^2

pi*r^2 = 3.07 * 10^-6

r = 9.89 * 10^-4 m = 0.989 mm

diameter = 2r = 1.98 mm

diameter of wire used is 1.98 mm
Answer #2

we have

R = \rho \frac{l}{A}

R = 14 ohms is resistance,

l = 43 m

\rho = 1.1 x 10-6 ohm m is resistivity

plugging in the valuesin the equation we get

A = 3.4 x 10-6 m2

we have

A = \pi r^{2}

that gives r = 1 mm

diameter d = 2r = 2mm